OpenStudy (anonymous):

cos(-870) Simplify...... I know to do such type of problems but this one has a - sign.... So does this minus go outside or should i keep it inside? Sin(-x) is - sinx right? Should i use that?

OpenStudy (anonymous):

Sure, use that.

OpenStudy (anonymous):

Don't use that..... Cosines are rather.... Resistant to negatives.

OpenStudy (anonymous):

Let me put it this way. When you're taking cosines (in radians or degrees) they're pretty much always positive within 90 degrees (or radians).

OpenStudy (anonymous):

So it becomes - cos(870) =-cos(720+150) =-cos(150) =-cos(180-30) =(-)(-cos30) =cos30?

OpenStudy (anonymous):

The second you move out of that, though, it gets a little weird. We're know talking about a unit circle.

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

So let's put it this way. If I have a circle, and I consider its center as an origin, everything on the left side of the circle is positive. Everything on the left side is negative. You with me?

OpenStudy (anonymous):

There was a typo there.... Sorry.... Cos(-x) is cosx right?

OpenStudy (anonymous):

Everything on the RIGHT side is positive. Everything on the LEFT is negative.

OpenStudy (anonymous):



OpenStudy (anonymous):

Precisely. You got it man.

OpenStudy (anonymous):

So the answer is cos30?

OpenStudy (anonymous):

Cos(-x) = Cos(x) There's proofs and whatnot (what I was going to show you) but you got it.

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

Forgetting the unit circle.